40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Oct 12 22:51:26 2016, Koji, Update, General, Clipping could not be improved by the PZTs - NEED ANOTHER VENT 
    Reply  Thu Oct 13 00:23:54 2016, ericq, Update, VAC, Pre-vent prep 
       Reply  Thu Oct 13 09:30:38 2016, Steve, Update, General, vent 79 is happening vent79precond.pngventing.png
          Reply  Thu Oct 13 16:48:34 2016, Steve, Update, VAC, vent 79 is completed vent_79.png
             Reply  Thu Oct 13 19:28:14 2016, gautam, Update, General, in-air alignment 
                Reply  Fri Oct 14 10:31:53 2016, gautam, Update, General, doors are off ITMY and BS/PRM chambers 
                   Reply  Fri Oct 14 18:33:55 2016, gautam, Update, General, AS clipping investigations 7x
                      Reply  Mon Oct 17 22:45:16 2016, gautam, Update, General, AS beam centered on all OMs 7x
                         Reply  Tue Oct 18 18:56:57 2016, gautam, Update, General, OM5 rotated to bypass OMC, green scatter is from window to PSL table OMCchamber.pdfgreenGhosts.JPGIMG_3322.JPGIMG_3326.JPG
                            Reply  Wed Oct 19 16:41:55 2016, gautam, Update, General, Heavy doors back on 8x
             Reply  Thu Oct 20 09:17:08 2016, Steve, Update, VAC, Pumpdown 80 has started readyPd80.pngbeforePd80.pngpd80at1h.png7d_atm.png
                Reply  Thu Oct 20 15:59:52 2016, Steve, Update, VAC, Pumpdown 80 has completed 7hPd80.png8hPd80.png9hPd80.png
                   Reply  Fri Oct 21 02:06:20 2016, gautam, Update, General, IFO recovery 
                      Reply  Mon Oct 24 11:39:13 2016, gautam, Update, General, ALS recovered ALSOutOfLoop20161024.pdfXendPDHOLTF20161024.pdf
                   Reply  Fri Oct 21 09:28:21 2016, Steve, Update, VAC, Vac Normal reached VacNormal.png
                      Reply  Wed Nov 23 11:32:13 2016, Steve, Update, VAC, TP3 drypump replaced again 
                         Reply  Fri Nov 10 08:52:21 2017, Steve, Update, VAC, TP3 drypump replaced 
                            Reply  Fri Feb 16 15:57:58 2018, Steve, Update, VAC, TP3 drypump replaced 
                Reply  Thu Jul 12 16:18:22 2018, Steve, Update, VAC, Vent preparations before_Vent.pngbefore_Vent_cond.png
                   Reply  Thu Jul 12 23:59:14 2018, gautam, Update, General, Vent objectives and prep 
                      Reply  Fri Jul 13 02:52:11 2018, gautam, Update, General, Vent objectives and prep preVentStatus.png
                         Reply  Fri Jul 13 17:13:45 2018, Koji, Update, General, Vent objectives and prep P_20180713_020603.jpg
                         Reply  Fri Jul 13 18:01:13 2018, gautam, Update, General, Low power MC 
Message ID: 12560     Entry time: Thu Oct 13 19:28:14 2016     In reply to: 12559     Reply to this: 12561
Author: gautam 
Type: Update 
Category: General 
Subject: in-air alignment 

I did the following today to prepare for taking the doors off tomorrow.

  • Locked MC at low power
    • low power autolocker used during the last vent isn't working so well now
    • so I manually locked the IMC - locks are holding for ~30 mins and MC transmission was maximized by tweaking MC1 and MC2 alignments. The transmission is now ~1150 which is what I remember it being from the last vent
    • I had to restart c1aux to run LSCoffsets
  • Aligned arms to green using bias sliders on IFO align
    • X green transmission is ~0.4 and Y green transmission is ~0.5 which is what I remember it being before this vent
  • Removed ND filters from end Transmon QPDs since there is so little light now
  • Locked Y arm, ran the dither
    • Some tip tilt beam walking was required before any flashes were seen
    • I had to tweak the LSC gain for this to work
    • TRY is ~0.3 - in the previous vent, in air low power locking yielded TRY of ~0.6 but the 50-50 BS that splits the light between the high gain PD and the QPD was removed back then so these numbers are consistent
  • Tried locking X arm
    • For some reason, I can't get the triggering to work well - the trigger in monitor channel (LSC-XARM_TRIG_IN) and LSC-TRX_OUT_DQ are not the same, should they not be?
    • Tried using both QPD and high gain PD to lock, no luck. I also checked the error signal for DC offsets and that the demod phase was okay
    • In any case, there are TRX flashes of ~0.3 as well, this plus the reasonable green transmission makes me think the X arm alignment is alright
  • All the oplev spots are on their QPDs in the +/- 100 range. I didn't bother centering them for now

I am leaving all shutters closed overnight.

So I think we are ready to take the doors off at 8am tomorrow morning, unless anyone thinks there are any further checks to be done first.


Vent objectives:

  1. Fix AS beam clipping issues (elog1, elog2)
  2. Look into the green scatter situation (elog)

Should we look to do anything else now? One thing that comes to mind is should we install ITM baffles? Or would this be more invasive than necessary for this vent?


Steve reported to me that he was unable to ssh into the control room machines from the laptops at the Xend and near the vacuum rack. The problem was with pianosa being frozen up. I did a manual reboot of pianosa and was able to ssh into it from both laptops just now.

ELOG V3.1.3-